What quantity of heat is needed to melt 120 g of ice at -20 o

c to water at 30 o

c? Specific 

heat capacity of water and ice are 4200 JKg-1K

-1 and 2100 JKg-1K

-1

respectively. And 

also to specific latent heat 0f ice as 3.36 x 105

JK-1


1
Expert's answer
2020-07-09T10:33:28-0400

We can find the quantity of heat from the formula:


"Q = Q_1+Q_2+Q_3,"

here, "Q_1 = mc_{ice} \Delta T_1" is the quantity of heat needed to change the temperature of the ice from "-20 \ ^{\circ}C" to "0 \ ^{\circ}C", "Q_2 = mL_f" is the quantity of heat needed to transform ice into the water at "0 \ ^{\circ}C", "Q_3 = mc_{water} \Delta T_3" is the quantity of heat needed to change the temperature of the water from "0 \ ^{\circ}C" to "30 \ ^{\circ}C", "m = 0.12 \ kg" is the mass of the substance, "c_{ice} = 2100 \ \dfrac{J}{kg \cdot \! ^{\circ}C}" is the specific heat capacity of ice, "c_{water} = 4200 \ \dfrac{J}{kg \cdot \! ^{\circ}C}" is the specific heat capacity of water, "L_f = 3.36 \cdot 10^5 \ \dfrac{J}{kg}" is the latent heat of fusion of ice.

Then, we get:


"Q = mc_{ice} \Delta T_1+mL_f+mc_{water} \Delta T_3,"

"Q = 0.12 \ kg \cdot(2100 \ \dfrac{J}{kg \cdot \! ^{\circ}C} \cdot 20 \ ^{\circ}C +3.36 \cdot 10^5 \ \dfrac{J}{kg}+4200 \ \dfrac{J}{kg \cdot \! ^{\circ}C} \cdot 30 \ ^{\circ}C)",


"Q = 60480 \ J."

Answer:

"Q = 60480 \ J".